OpenStudy (anonymous):

Would it be easier to factor 12t^2-17t^-1-5=0 or to use the quadratic formula?

OpenStudy (ash2326):

Is this your equation? \[12t^2-17t^{-1}-5=0\]

OpenStudy (anonymous):

i dont think you can use the quadratic formula in this case, as we have -ve power in this case.

OpenStudy (anonymous):

\[12t^2-17t^{-1}-5=0\]This one won't factor that well. The problem is probably\[12t^2-17t-5=0\]

OpenStudy (anonymous):

If it is, you can use the ac-method to factor it. Do you know how to do that?

OpenStudy (anonymous):

Yes, but that's not the method my teacher wants me to use. Thank you.

OpenStudy (anonymous):

what is the problem exactly. is it quadratic or not?

OpenStudy (anonymous):

No. it's not quadratic.

OpenStudy (anonymous):

The quadratic formula only works on quadratic equations.

OpenStudy (anonymous):

there must have been a mistake with this problem. anyways, the equation can be reduced to cubic form.the equation has a sign change between 1 and 2. so it must have a root between these two numbers. i found it to be 1.246. the other two roots can be found by reducing this cubic equation into quadratic equation and they come out to be imaginary.